# JEL classification: C33, C38, C55

### v1

@@ A Mathematical Proofs
$|\mathcal{N}|\leq\left(\begin{array}[]{c}N\\ 2s_{1}\end{array}\right)(1+\frac{2}{\delta})^{2s_{1}}\leq\left(\frac{eN}{2s_{1}}\right)^{2s_{1}}\left(1+\frac{2}{\delta}\right)^{2s_{1}},$
@@
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|\leq(1-2\delta)^{-1}\sup_{{\mathbf{v}}\in\mathcal{N}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|.$
We choose $\delta=1/4$, by the Bonferroni-type inequality and Lemma A.6,
@@
$\leq$ $C\left(\frac{eN}{2s_{1}}\right)^{2s_{1}}9^{2s_{1}}\exp\left(-C\min\left({NT^{2}t^{2}},NTt\right)\right).$ (IA.15)
Therefore, we can obtain from (A) that
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|\leq C\max\{\sqrt{\frac{s_{1}\log(N)}{NT^{2}}},\frac{s_{1}\log(N)}{NT}\}\leq C\sqrt{\frac{s_{1}\log(N)}{NT^{2}}}+C\frac{s_{1}\log(N)}{NT}.$
Combining (IA.14) and the above inequality, we have
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}{\mathbf{S}}_{e}{\mathbf{v}}|\leq C\left(\sqrt{\frac{s_{1}\log(N)}{NT^{2}}}+\frac{s_{1}\log(N)}{NT}+\frac{1}{T}\right).$ (IA.16)
@@
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}\boldsymbol{\Sigma}_{fe}{\mathbf{v}}|\leq C\left(\sqrt{\frac{s_{1}\log(N)}{NT}}\right).$ (IA.17)
By Assumption 4 that $\lambda_{1}-\lambda_{2}>c>0$, Theorem 1 follows from Lemma A.4, (IA.1), and the above two inequalities. This completes the proof. $\Box$
@@
$C_{NT}=\sqrt{\frac{s_{0}\log(N)}{NT}}+\frac{s_{0}\log(N)}{NT}+\frac{1}{T},$
@@
$\widehat{\boldsymbol{\Lambda}}^{\prime}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{X}}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{F}}\boldsymbol{\Lambda}^{\prime}+(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}.$ (IA.18)
It follows from (IA.18) that
$\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i}=(\frac{\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}}{T})^{-1}\frac{\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}}{T},$ (IA.19)
@@
$\max_{1\leq i\leq N}|\frac{\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}}{T}|=O_{p}(\sqrt{\frac{s^{*}}{NT}}\log(N)+\frac{s^{*}\log(N)^{3/2}}{NT}+\sqrt{\frac{\log(N)}{T}}).$
If $s^{*}$ is finite, then it reduces to
@@
Note that $\frac{\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}}{T}\rightarrow_{p}{\mathbf{I}}_{r}$, by Assumptions 56 and the results in Theorem 2, we can show that
$\sqrt{T}(\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i})\longrightarrow_{d}N(0,\sigma_{i}^{2}{\mathbf{I}}_{r}),$ (IA.20)
@@
Proof. We only show it for $r=1$ since the proof is similar for $r>1$. Suppose ${\mathbf{F}}^{s}$ is the normalized true factor matrix such that ${\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s}/T={\mathbf{I}}_{r}$. By (14), we have
$R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})=\frac{1}{N_{2}T}\|\widetilde{\mathbf{X}}_{2}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2}\|_{F}^{2}-\frac{1}{N_{2}T}\|\widetilde{\mathbf{X}}_{2}-\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2}\|_{F}^{2}.$ (IA.21)
@@
$R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})=\frac{1}{N_{2}T}\mbox{tr}\{(\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}})\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\}.$ (IA.22)
@@
$\|\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\|_{2}=\|\frac{1}{\sqrt{T}}({\mathbf{F}}^{s}-\widetilde{\mathbf{F}}^{s})\frac{{\mathbf{F}}^{s}{{}^{\prime}}}{\sqrt{T}}+\frac{\widetilde{\mathbf{F}}^{s}}{\sqrt{T}}\frac{({\mathbf{F}}^{s}-\widetilde{\mathbf{F}}^{s})^{\prime}}{\sqrt{T}}\|_{2}=O_{p}(C_{N_{1}T}),$ (IA.23)
@@
it follows from (IA.22) that
@@
Proof. By a similar argument as that in (IA.22), we can obtain that
$R(s,{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})=\frac{1}{N_{2}T}\mbox{tr}\{[{\mathbf{F}}^{0}({\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{F}}^{0})^{-1}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}]\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\},$ (IA.24)
@@
$\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}={\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}+{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}{{}^{\prime}}{\mathbf{E}}_{2}^{\prime}+{\mathbf{E}}_{2}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}+{\mathbf{E}}_{2}{\mathbf{E}}_{2}^{\prime},$ (IA.25)
where $\boldsymbol{\Lambda}_{2}$ and ${\mathbf{E}}_{2}$ are the loading matrix and idiosyncratic term of $\widetilde{\mathbf{X}}_{2}$. Then,
@@
$=$ $\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}\boldsymbol{\Lambda}_{2}^{\prime}{\mathbf{F}}^{0}-{\mathbf{F}}^{s}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}$
@@
$I=\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}\boldsymbol{\Lambda}_{2}^{\prime}{\mathbf{F}}^{0}-{\mathbf{F}}^{s}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}\},$
@@
Since each column in ${\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}$ is a hat matrix and ${\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{G}}^{\prime}$ is a positive semi-definite matrix, then there exists a constant $\tau_{s}>0$ such that $I\geq\tau_{s}N_{2}T$. Then Lemma A.9 follows from this result and (IA.24)–(A). This completes the proof. $\Box$
@@
$\leq$ $2\max_{s_{0}\leq s\leq smax}|R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})|.$ (IA.27)
@@
$|R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})|=O_{p}(C_{N_{1}T}^{2}).$ (IA.28)
@@
$=$ $\frac{1}{N_{2}T}\mbox{tr}({\mathbf{L}}).$ (IA.29)
By (IA.25), we can show that
@@
$-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{E}}_{2}{\mathbf{E}}_{2}^{\prime}.$ (IA.30)
@@
$=$ $\mbox{tr}\{({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}\}$ (IA.31)
@@
$\frac{1}{N_{2}T}\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}\}=O_{p}(C_{N_{1}T}^{2}).$ (IA.32)
By a similar argument for the remaining terms in (A), we can show that
@@
which implies (IA.28). This completes the proof. $\Box$
@@
$PC(k)-PC(s_{0})=R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},\widetilde{\mathbf{F}}^{0})-r(s_{0}-s)g(N_{1},T),$
@@
$P(PC(s)-PC(s_{0})<0)=P(R(s_{0},\widetilde{\mathbf{F}}^{0})-R(s,\widetilde{\mathbf{F}}^{s})>r(s-s_{0})g(N_{1},T)).$
@@
B Description of the Identified Sparse Factors

### v2

@@ B Mathematical Proofs
$|\mathcal{N}|\leq\left(\begin{array}[]{c}T\\ 2s_{1}\end{array}\right)(1+\frac{2}{\delta})^{2s_{1}}\leq\left(\frac{eT}{2s_{1}}\right)^{2s_{1}}\left(1+\frac{2}{\delta}\right)^{2s_{1}},$
@@
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|\leq(1-2\delta)^{-1}\sup_{{\mathbf{v}}\in\mathcal{N}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|,$
because the left-hand-side (LHS) satisfies $LHS\leq 2\delta*LHS+\sup_{{\mathbf{v}}\in\mathcal{N}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|$. We choose $\delta=1/4$, by the Bonferroni-type inequality and Lemma A.6,
@@
$\leq$ $C\left(\frac{eT}{2s_{1}}\right)^{2s_{1}}9^{2s_{1}}\exp\left(-C\min\left({NT^{2}t^{2}},NTt\right)\right).$ (S.15)
Therefore, we can obtain from (B) that
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|\leq C\max\{\sqrt{\frac{s_{1}\log(T)}{NT^{2}}},\frac{s_{1}\log(T)}{NT}\}\leq C\sqrt{\frac{s_{1}\log(T)}{NT^{2}}}+C\frac{s_{1}\log(T)}{NT}.$
Combining (S.14) and the above inequality, we have
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}{\mathbf{S}}_{e}{\mathbf{v}}|\leq C\left(\sqrt{\frac{s_{1}\log(T)}{NT^{2}}}+\frac{s_{1}\log(T)}{NT}+\frac{1}{T}\right).$ (S.16)
@@
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}\boldsymbol{\Sigma}_{fe}{\mathbf{v}}|\leq C\left(\sqrt{\frac{s_{1}\log(T)}{NT}}\right).$ (S.17)
By Assumption 4 that $\lambda_{1}-\lambda_{2}>c>0$, Theorem 1 follows from Lemma A.4, (S.1), and the above two inequalities. This completes the proof. $\Box$
@@
$C_{NT}=\sqrt{\frac{s_{0}\log(T)}{NT}}+\frac{s_{0}\log(T)}{NT}+\frac{1}{T},$
@@
$\widehat{\boldsymbol{\Lambda}}^{\prime}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{X}}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{F}}\boldsymbol{\Lambda}^{\prime}+(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}.$ (S.18)
It follows from (S.18) that
$\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i}=(\frac{\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}}{T})^{-1}\frac{\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}}{T},$ (S.19)
@@
$\max_{1\leq i\leq N}|\frac{\widehat{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}}{T}|$ $\leq\max_{1\leq i\leq N}|\frac{1}{T}(\widehat{\mathbf{F}}-{\mathbf{F}})^{\prime}{\mathbf{e}}_{i}|+\max_{1\leq i\leq N}|\frac{1}{T}{\mathbf{F}}^{\prime}{\mathbf{e}}_{i}|$
$\leq=O_{p}(\sqrt{\frac{s^{*}{{}^{2}}\log(N)}{NT^{2}}}\log(T)+\sqrt{\frac{\log(N)}{T}}).$
Since $s^{*}=T^{\alpha}$ where $\alpha=\max(\alpha_{1},...,\alpha_{r})$ and $\alpha_{i}$ is defined in Assumption 3, if $\sqrt{\frac{1}{NT^{1-2\alpha}}}\log(T)=o(1)$, then it reduces to
@@
Let $\frac{\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}}{T}\rightarrow_{p}{\mathbf{Q}}$ and $\boldsymbol{\Gamma}_{i}=\lim_{T\rightarrow\infty}\textnormal{Var}(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}{\mathbf{f}}_{t}e_{i,t})$, it is straightforward that
$\sqrt{T}(\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i})\longrightarrow_{d}N(0,{\mathbf{Q}}^{-1}\boldsymbol{\Gamma}_{i}{\mathbf{Q}}^{-1}).$
Note that $\frac{\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}}{T}\rightarrow_{p}{\mathbf{I}}_{r}$ due to the average consistency of each estimated factor process and Assumption 2, by Assumptions 56 and the results in Theorem 2, we can show that
$\sqrt{T}(\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i})\longrightarrow_{d}N(0,\sigma_{i}^{2}{\mathbf{I}}_{r}),$ (S.20)
@@
Proof. We only show it for $r=1$ since the proof is similar for $r>1$. Suppose the true factor ${\mathbf{F}}^{s}$ is the normalized true factor matrix such that ${\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s}/T={\mathbf{I}}_{r}$, which is similarly assumed in Bai and Ng (2013). By (14), we have
$R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})=\frac{1}{N_{2}T}\|\widetilde{\mathbf{X}}_{2}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2}\|_{F}^{2}-\frac{1}{N_{2}T}\|\widetilde{\mathbf{X}}_{2}-\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{X}}_{2}\|_{F}^{2}.$ (S.21)
@@
$R(s,\widetilde{\mathbf{F}}^{s})-R(s,{\mathbf{F}}^{s})=\frac{1}{N_{2}T}\mbox{tr}\{(\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}})\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\}.$ (S.22)
@@
$\|\frac{1}{T}{\mathbf{F}}^{s}{\mathbf{F}}^{s}{{}^{\prime}}-\frac{1}{T}\widetilde{\mathbf{F}}^{s}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\|_{2}=\|\frac{1}{\sqrt{T}}({\mathbf{F}}^{s}-\widetilde{\mathbf{F}}^{s})\frac{{\mathbf{F}}^{s}{{}^{\prime}}}{\sqrt{T}}+\frac{\widetilde{\mathbf{F}}^{s}}{\sqrt{T}}\frac{({\mathbf{F}}^{s}-\widetilde{\mathbf{F}}^{s})^{\prime}}{\sqrt{T}}\|_{2}=O_{p}(C_{N_{1}T}),$ (S.23)
@@
it follows from (S.22) that
@@
Proof. By a similar argument as that in (S.22), we can obtain that
$R(s,{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})=\frac{1}{N_{2}T}\mbox{tr}\{[{\mathbf{F}}^{0}({\mathbf{F}}^{0}{{}^{\prime}}{\mathbf{F}}^{0})^{-1}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}]\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}\},$ (S.24)
@@
$\widetilde{\mathbf{X}}_{2}\widetilde{\mathbf{X}}_{2}^{\prime}={\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}+{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}{{}^{\prime}}{\mathbf{E}}_{2}^{\prime}+{\mathbf{E}}_{2}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}+{\mathbf{E}}_{2}{\mathbf{E}}_{2}^{\prime},$ (S.25)
where $\boldsymbol{\Lambda}_{2}$ and ${\mathbf{E}}_{2}$ are the loading matrix and idiosyncratic term of $\widetilde{\mathbf{X}}_{2}$, respectively. Then,
@@
$=$ $\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}$
@@
$I=\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-{\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}\},$
@@
Since each column in ${\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}$ is a hat matrix and ${\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{G}}^{\prime}$ is a positive semi-definite matrix, then there exists a constant $\tau_{s}>0$ such that $I\geq\tau_{s}N_{2}T$. Then Lemma A.9 follows from this result and (S.24)–(B). This completes the proof. $\Box$
@@
$\leq$ $2\max_{s_{0}\leq s\leq smax}|R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})|.$ (S.27)
@@
$|R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},{\mathbf{F}}^{0})|=O_{p}(C_{N_{1}T}^{2}).$ (S.28)
@@
$=$ $\frac{1}{N_{2}T}\mbox{tr}({\mathbf{L}}).$ (S.29)
By (S.25), we can show that
@@
$-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{E}}_{2}{\mathbf{E}}_{2}^{\prime}.$ (S.30)
@@
$=$ $\mbox{tr}\{({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}\}$ (S.31)
@@
$\frac{1}{N_{2}T}\mbox{tr}\{{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{0}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{F}}^{0}{{}^{\prime}}\}=O_{p}(C_{N_{1}T}^{2}).$ (S.32)
By a similar argument for the remaining terms in (B), we can show that
@@
which implies (S.28). This completes the proof. $\Box$
@@
$PC(s)-PC(s_{0})=R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},\widetilde{\mathbf{F}}^{0})-r\frac{(s_{0}-s)}{T^{\alpha}}g(N_{1},T),$
@@
$P(PC(s)-PC(s_{0})<0)=P(R(s_{0},\widetilde{\mathbf{F}}^{0})-R(s,\widetilde{\mathbf{F}}^{s})>r\frac{(s-s_{0})}{T^{\alpha}}g(N_{1},T)).$
@@
C Description of the Identified Sparse Factors